Required information

The Foundational 15 [LO2-1, LO2-2, LO2-3, LO2-4]

[The following information applies to the questions displayed below.]

Sweeten Company had no jobs in progress at the beginning of March and no beginning inventories. The company has two manufacturing departments—Molding and Fabrication. It started, completed, and sold only two jobs during March—Job P and Job Q. The following additional information is available for the company as a whole and for Jobs P and Q (all data and questions relate to the month of March):

Molding Fabrication Total
Estimated total machine-hours used 2,500 1,500 4,000
Estimated total fixed manufacturing overhead $ 12,750 $ 16,650 $ 29,400
Estimated variable manufacturing overhead per machine-hour $ 2.50 $ 3.30
Job P Job Q
Direct materials $ 24,000 $ 13,500
Direct labor cost $ 29,800 $ 11,900
Actual machine-hours used:
Molding 2,800 1,900
Fabrication 1,700 2,000
Total 4,500 3,900

Sweeten Company had no underapplied or overapplied manufacturing overhead costs during the month.

Required:

For questions 1-9, assume that Sweeten Company uses departmental predetermined overhead rates with machine-hours as the allocation base in both departments and Job P included 20 units and Job Q included 30 units. For questions 10-15, assume that the company uses a plantwide predetermined overhead rate with machine-hours as the allocation base.

Foundational 2-8 (Algo)

8. Assume that Sweeten Company used cost-plus pricing (and a markup percentage of 80% of total manufacturing cost) to establish selling prices for all of its jobs. What selling price would the company have established for Jobs P and Q? What are the selling prices for both jobs when stated on a per unit basis? (Do not round intermediate calculations. Round your final answers to nearest whole dollar.)

9. What was Sweeten Company’s cost of goods sold for March? (Do not round intermediate calculations.)

12. If Job P included 20 units, what was its unit product cost? (Do not round intermediate calculations. Round your final answer to nearest whole dollar.)

13. If Job Q included 30 units, what was its unit product cost? (Do not round intermediate calculations. Round your final answer to nearest whole dollar.)

14. Assume that Sweeten Company used cost-plus pricing (and a markup percentage of 80% of total manufacturing cost) to establish selling prices for all of its jobs. What selling price would the company have established for Jobs P and Q? What are the selling prices for both jobs when stated on a per unit basis? (Do not round intermediate calculations. Round your final answers to nearest whole dollar.)

15. What was Sweeten Company’s cost of goods sold for March? (Do not round intermediate calculations.)
Answer #1

Molding overhead rate = (12750/2500+2.5) = 7.60 per hour

Fabrication overhead rate = (16650/1500+3.3) = 14.40 per hour

8) Calculate following

Job P Job Q
Direct material 24000 13500
Direct labor 29800 11900
Molding overhead 2800*7.6 = 21280 14440
Fabrication overhead 1700*14.40 = 24480 28800
Total 99560 68640
Selling price for the job 179208 68640*1.8 = 123552
Unit price 8960 4118

9) Cost of goods sold = 99560+68640 = 168200

Plantwide rate = (2500*2.5+1500*3.3+29400)/4000 = 10.15 per hour

12) Unit product cost

Job P
Direct material 24000
Direct labor 29800
Overhead 4500*10.15 = 45675
Total 99475
Unit cost 4974

13) Unit product cost

Job Q
Direct material 13500
Direct labor 11900
Overhead 3900*10.15 = 39585
Total 64985
Unit cost 2166